# Underdetermined Fourier Extensions for Surface Partial Differential Equations

### v2

@@ 3.5. Analysis for Hermite-Birkhoff Problem
$\leq$ $\ A_{k,q}\max_{\left|\beta\right|\leq q}\left(\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}\tilde{b}_{n}\left\|\partial^{\beta}\mathcal{G}^{\left(j\right)}\phi_{n}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}+\left\|\partial^{\beta}g^{\left(j\right)}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}\right)$
[…]
Then there exist constants $\tilde{A}_{k,q}>0$ so that
$\left\|\left(\mathcal{G}^{\left(j\right)}\tilde{u}-g\right)\circ\sigma^{\left(j\right)}_{k}\right\|_{H^{q}\left(U_{k}^{\left(j\right)}\right)}\leq\tilde{A}_{k,q}\left(\tilde{C}_{j,q}\left\|\tilde{b}\right\|_{\ell^{2}}+\max_{\left|\beta\right|\leq q}\left(\left\|\partial^{\beta}g^{\left(j\right)}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}\right)\right).$
@@
Once again using Corollary 7, we have that there exist constants $\tilde{B}_{k,q}>0$ so that, for small enough $h_{\text{max}}$,
@@
$\left\|\left(\mathcal{G}^{\left(j\right)}\tilde{u}-g\right)\circ\sigma^{\left(j\right)}_{k}\right\|_{L^{\infty}\left(U_{k}^{\left(j\right)}\right)}\leq\tilde{B}_{k,q}h_{\text{max}}^{q-\frac{m_{j}}{2}}\left(\tilde{C}_{j,q}\left\|u\right\|_{d}+\max_{\left|\beta\right|\leq q}\left(\left\|\partial^{\beta}g^{\left(j\right)}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}\right)\right).$
@@
Note that Corollary 7 is applied to $U^{\left(j\right)}_{k}$, not $S^{\left(j\right)}_{k}$, so we are implicitly relying on the boundedness of the inverse metric tensor of $\sigma^{\left(j\right)}_{k}$ so that the fill distance on $U^{\left(j\right)}_{k}$ is at most a constant multiple of the fill distance on $S^{\left(j\right)}_{k}$.
Finally, letting $\tilde{B}_{q}=\max\left\{\tilde{B}_{k,q}\right\}_{k=1}^{M}$, we have the desired result:
$\left\|\mathcal{G}^{\left(j\right)}\tilde{u}-g^{\left(j\right)}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}\leq\tilde{B}_{q}h_{\text{max}}^{q-\frac{m_{j}}{2}}\left(\tilde{C}_{j,q}\left\|u\right\|_{d}+\max_{\left|\beta\right|\leq q}\left(\left\|\partial^{\beta}g^{\left(j\right)}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}\right)\right).$
@@ 3.6. Finite Number of Basis Functions
We now truncate our Fourier series to a finite basis of $N_{b}$ terms. Let $\mathcal{L}_{N_{b}}:=\mathcal{L}_{\text{HB}}\biggr{|}_{\mathbb{C}^{N_{b}}}:\mathbb{C}^{N_{b}}\to\mathbb{C}^{\tilde{N}}$, where $\mathcal{L}_{\text{HB}}$ is as defined in the proof of Proposition 8. Let $\boldsymbol{f}:=\left(f^{\left(j\right)}\left(\boldsymbol{x}_{k}^{\left(j\right)}\right)\right)$, $\boldsymbol{g}:=\left(g^{\left(j\right)}\left(\boldsymbol{y}_{k}^{\left(j\right)}\right)\right)$ be vectorized forms of all Hermite-Birkhoff data from (3.4) (column vectors using all values of $j$ and $k$). The problem then becomes:
[…]
(3.9) subject to: $\mathcal{L}_{N_{b}}\boldsymbol{b}=\text{proj}_{\mathcal{R}\left(\mathcal{L}_{N_{b}}\right)}\left(\begin{array}[]{c}\boldsymbol{f}\\ \boldsymbol{g}\end{array}\right).$
@@
$\text{proj}_{\mathcal{R}\left(\mathcal{L}_{N_{b}}\right)}\left(\begin{array}[]{c}\boldsymbol{f}\\ \boldsymbol{g}\end{array}\right)=\left(\begin{array}[]{c}\boldsymbol{f}\\ \boldsymbol{g}\end{array}\right).$
@@
Assume there is a constant $B_{p}$ such that for all large enough $n$, $\left\|\mathcal{F}^{\left(j\right)}\phi_{n}\right\|_{L^{\infty}\left(\Omega^{\left(j\right)}\right)}\leq B_{p}\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{p}$ for each $j\in\left\{1,2,\ldots,N_{\Omega}\right\}$ and $\left\|\mathcal{G}^{\left(j\right)}\phi_{n}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}\leq B_{p}\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{p}$ for each $j\in\left\{1,2,\ldots,N_{S}\right\}$. Also assume $\left\|\boldsymbol{\omega}\right\|_{2}^{p}d^{-\frac{1}{2}}\in\ell^{2}$. Then, for a fixed number of total Hermite-Birkhoff interpolation conditions $\tilde{N}$, there exists a constant $A_{\tilde{N}}>0$ such that for sufficiently large $N_{b}$,
@@
The approach of the proof is to note that if an interpolant exists for finite $N_{b}$, it is also feasible for the $\ell^{2}$ problem, while the truncated solution to the $\ell^{2}$ problem will “nearly” be an interpolant and will be close to a feasible solution for the finite basis problem. Along with orthogonality of the minimizer to the null space, this turns out to be enough to show convergence proportional to $\left(\sum_{n=N_{b}+1}^{\infty}\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{2p}d_{n}^{-1}\right)^{\frac{1}{2}}$ of the truncated basis solution to the $\ell^{2}$ problem. The idea here is to show that adding terms to our Fourier basis recovers the solution to the full $\ell^{2}$ problem in the limit $N_{b}\to\infty$.
@@
$\boldsymbol{V}_{N_{b}}\tilde{\boldsymbol{b}}$ $=\left(\begin{array}[]{c}\boldsymbol{f}\\ \boldsymbol{g}\end{array}\right),$
@@
$\tilde{\boldsymbol{b}}=\boldsymbol{V}_{N_{b}}^{+}\left(\begin{array}[]{cc}\boldsymbol{f}\\ \boldsymbol{g}\end{array}\right),$
@@
This holds for any $\boldsymbol{f},\boldsymbol{g}$, so if we define $M_{\tilde{N}}=\left\|\boldsymbol{V}_{N_{b,\text{interp}}}^{+}\right\|_{2}$ , then as long as $N_{b}\geq N_{b,\text{interp}}$,
@@
$\varepsilon_{\text{trunc}}:=\left\|\left(\begin{array}[]{c}\boldsymbol{f}\\ \boldsymbol{g}\end{array}\right)-\boldsymbol{V}_{N_{b}}\tilde{\boldsymbol{b}}_{\infty}^{\left(N_{b}\right)}\right\|_{2},$
@@
$\boldsymbol{V}_{N_{b}}\left(\tilde{\boldsymbol{b}}_{\infty}^{\left(N_{b}\right)}+\boldsymbol{V}_{N_{b}}^{+}\left(\left(\begin{array}[]{c}\boldsymbol{f}\\ \boldsymbol{g}\end{array}\right)-\boldsymbol{V}_{N_{b}}\tilde{\boldsymbol{b}}_{\infty}^{\left(N_{b}\right)}\right)\right)=\left(\begin{array}[]{c}\boldsymbol{f}\\ \boldsymbol{g}\end{array}\right).$
@@
(3.12) $\left\|\tilde{\boldsymbol{b}}_{\infty}^{\left(N_{b}\right)}+\boldsymbol{V}_{N_{b}}^{+}\left(\left(\begin{array}[]{c}\boldsymbol{f}\\ \boldsymbol{g}\end{array}\right)-\boldsymbol{V}_{N_{b}}\tilde{\boldsymbol{b}}_{\infty}^{\left(N_{b}\right)}\right)\right\|_{2}$ $\geq\left\|\tilde{\boldsymbol{b}}\right\|_{2}$
@@
$\varepsilon_{\text{trunc}}^{2}$ $=\sum_{j=1}^{N_{\Omega}}\sum_{k=1}^{\tilde{N}_{j}}\left|f^{\left(j\right)}\left(\boldsymbol{x}^{\left(j\right)}_{k}\right)-\sum_{n=1}^{N_{b}}d_{n}^{-\frac{1}{2}}\tilde{b}_{\infty,n}\mathcal{F}^{\left(j\right)}\phi_{n}\left(\boldsymbol{x}^{\left(j\right)}_{k}\right)\right|^{2}$
$\quad+\sum_{j=1}^{N_{S}}\sum_{k=1}^{\tilde{M}_{j}}\left|g^{\left(j\right)}\left(\boldsymbol{y}^{\left(j\right)}_{k}\right)-\sum_{n=1}^{N_{b}}d_{n}^{-\frac{1}{2}}\tilde{b}_{\infty,n}\mathcal{G}^{\left(j\right)}\phi_{n}\left(\boldsymbol{y}^{\left(j\right)}_{k}\right)\right|^{2}$
$=\sum_{j=1}^{N_{\Omega}}\sum_{k=1}^{\tilde{N}_{j}}\left|\sum_{n=N_{b}+1}^{\infty}d_{n}^{-\frac{1}{2}}\tilde{b}_{\infty,n}\mathcal{F}^{\left(j\right)}\phi_{n}\left(\boldsymbol{x}^{\left(j\right)}_{k}\right)\right|^{2}$
$\quad+\sum_{j=1}^{N_{S}}\sum_{k=1}^{\tilde{M}_{j}}\left|\sum_{n=N_{b}+1}^{\infty}d_{n}^{-\frac{1}{2}}\tilde{b}_{\infty,n}\mathcal{G}^{\left(j\right)}\phi_{n}\left(\boldsymbol{y}^{\left(j\right)}_{k}\right)\right|^{2}$
[…]
where the last line holds for large enough $N_{b}$ by assumption on $\mathcal{F}^{\left(j\right)},\mathcal{G}^{\left(j\right)}$. Finally, from (3.14),
@@
In 1D, an interpolant for a combination of function or derivative interpolation conditions (a Hermite-Birkhoff interpolant) exists under certain conditions when $N_{b}=\tilde{N}$ [johns75], but this is not necessarily the case in dimensions greater than one, again due to the Mairhuber-Curtis Theorem [curti59, mairh56]. An interpolant does, however, always exist with some finite number of basis functions.
In particular, it can be quickly shown that $N_{b,\text{interp}}\leq\left(\left(p+1\right)\tilde{N}_{\text{distinct}}+1\right)^{m}$, where $\tilde{N}_{\text{distinct}}\leq\tilde{N}$ is the number of distinct points in the box $\Omega\subset\mathbb{R}^{m}$. $\tilde{N}_{\text{distinct}}$ can be less than $\tilde{N}$ if one or more points have two or more interpolation conditions imposed. For $m=1$, it has been known for some time [johns75, Thm. 1] that we can always construct a trigonometric polynomial with at most $\left(p+1\right)\tilde{N}_{\text{distinct}}+1$ terms that is zero and has all derivatives of up to order $p$ equal to zero on up to $\tilde{N}_{\text{distinct}}$ distinct points of our choosing, except for one chosen point where either the function or one of its derivatives are equal to one. For $m>1$, we can take products of these 1D functions to get a trigonometric polynomial with at most $\left(\left(p+1\right)\tilde{N}_{\text{distinct}}+1\right)^{m}$ terms that is zero and has derivatives equal to zero on a tensor grid of $\tilde{N}_{\text{distinct}}^{m}$ points (that our desired $\tilde{N}_{\text{distinct}}$ points are a subset of), except for one point and derivative of interest. A linear combination of these functions is a feasible interpolant. Experimentally, however, we typically observe $N_{b,\text{interp}}$ to simply be the total number of interpolation conditions ($\tilde{N}$) or not much larger (less than $2\tilde{N}$).
Theorem 2.2 of [chand18] provides an upper bound on $N_{b,\text{interp}}$ that depends on point separation and will often be less than the simple bound presented above. Note that increasing $N_{b}$ well beyond $N_{b,\text{interp}}$ can improve the error, in accordance with Proposition 11.

### v3

@@ 3.5. Analysis for Hermite-Birkhoff Problem
$\leq$ $\ A_{k,q}\max_{\left|\beta\right|\leq q}\left(\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}\left|\tilde{b}_{n}-b_{n}\right|\left\|\partial^{\beta}\mathcal{F}^{\left(j\right)}\phi_{n}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}\right)$
$\leq$ $\ A_{k,q}\tilde{C}_{j,q}\left\|\tilde{b}-b\right\|_{\ell^{2}}\text{, by Cauchy-Schwarz}.$
Now we note that $b$ is feasible for (3.5), so $\tilde{b}-b\in\mathcal{N}\left(\mathcal{L}_{\text{HB}}\right)$ in the notation of Subsection 3.3, so $\tilde{b}-b\perp\tilde{b}$ by (3.3). Importantly, $\left\|\tilde{b}-b\right\|_{\ell_{2}}=\sqrt{\left\|b\right\|_{\ell_{2}}^{2}-\left\|\tilde{b}\right\|^{2}_{\ell_{2}}}\leq\left\|b\right\|_{\ell_{2}}$. Then there exist constants $\tilde{A}_{j,k,q}>0$ so that
$\left\|\left(\mathcal{F}^{\left(j\right)}\tilde{u}-f^{\left(j\right)}\right)\circ\sigma^{\left(j\right)}_{k}\right\|_{H^{q}\left(U_{k}^{\left(j\right)}\right)}\leq\tilde{A}_{j,k,q}\left\|b\right\|_{\ell^{2}}.$
@@
Once again, using Corollary 6, we have that there exist constants $\tilde{B}_{k,q,\left|\alpha\right|,j}>0$ so that, for small enough $h_{\text{max}}$,
@@
$\left\|\partial^{\alpha}\left(\left(\mathcal{F}^{\left(j\right)}\tilde{u}-f^{\left(j\right)}\right)\circ\sigma_{k}^{\left(j\right)}\right)\right\|_{L^{\infty}\left(U_{k}^{\left(j\right)}\right)}\leq\tilde{B}_{j,k,q,\left|\alpha\right|}h_{\text{max}}^{q-\frac{m_{j}}{2}-\left|\alpha\right|}\left\|u\right\|_{d}.$
@@
Note that Corollary 6 is applied to $U^{\left(j\right)}_{k}$, not $S^{\left(j\right)}_{k}$, so we are implicitly relying on the boundedness of the inverse metric tensor of $\sigma^{\left(j\right)}_{k}$ so that the fill distance on $U^{\left(j\right)}_{k}$ is at most a constant multiple of the fill distance on $S^{\left(j\right)}_{k}$. ∎
In particular, notice that we can let $\tilde{B}_{j,q}=\max\left\{\tilde{B}_{j,k,q,0}\right\}_{k=1}^{M}$, then
$\left\|\mathcal{F}^{\left(j\right)}\tilde{u}-f^{\left(j\right)}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}\leq\tilde{B}_{j,q}h_{\text{max}}^{q-\frac{m_{j}}{2}}\left\|u\right\|_{d}.$
We note that the $F^{\left(j,q\right)}d^{-\frac{1}{2}}\in\ell^{2}$ condition tells us how quickly our choice of $d^{-\frac{1}{2}}$ must decay. If $\mathcal{F}^{\left(j\right)}$ is of order $p$ and has bounded coefficient functions with $q$ bounded derivatives, then $F^{\left(j,q\right)}$ will grow as $\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{p+q}$ as $n\to\infty$.
@@ 3.6. Finite Number of Basis Functions
We now truncate our Fourier series to a finite basis of $N_{b}$ terms. Let $\mathcal{L}_{N_{b}}:=\mathcal{L}_{\text{HB}}\biggr{|}_{\mathbb{C}^{N_{b}}}:\mathbb{C}^{N_{b}}\to\mathbb{C}^{\tilde{N}}$, where $\mathcal{L}_{\text{HB}}$ is as defined in the proof of Proposition 7. Let $\boldsymbol{f}:=\left(f^{\left(j\right)}\left(\boldsymbol{x}_{k}^{\left(j\right)}\right)\right)$ be a vectorized form of all Hermite-Birkhoff data from (3.5) (a column vector using all values of $j$ and $k$). The problem then becomes:
(3.6) $\text{minimize, for }\boldsymbol{b}\in\mathbb{C}^{N_{b}}:$ $\left\|\boldsymbol{b}\right\|_{2}$
subject to: $\mathcal{L}_{N_{b}}\boldsymbol{b}=\text{proj}_{\mathcal{R}\left(\mathcal{L}_{N_{b}}\right)}\boldsymbol{f}.$
@@
$\text{proj}_{\mathcal{R}\left(\mathcal{L}_{N_{b}}\right)}\boldsymbol{f}=\boldsymbol{f}.$
@@
Assume there is a constant $B_{p}$ such that for all large enough $n$, $\left\|\mathcal{F}^{\left(j\right)}\phi_{n}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}\leq B_{p}\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{p}$ for each $j\in\left\{1,2,\ldots,N_{S}\right\}$. Also assume $\left\|\boldsymbol{\omega}\right\|_{2}^{p}d^{-\frac{1}{2}}=\left\{\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{p}d_{n}^{-\frac{1}{2}}\right\}_{n=1}^{\infty}\in\ell^{2}$. Then, for a fixed number of total Hermite-Birkhoff interpolation conditions $\tilde{N}$, there exists a constant $A_{\tilde{N}}>0$ such that for sufficiently large $N_{b}$,
@@
The approach of the proof is to note that if an interpolant exists for finite $N_{b}$, it is also feasible for the $\ell^{2}$ problem, while the truncated solution to the $\ell^{2}$ problem will “nearly” be an interpolant and will be close to a feasible solution for the finite basis problem. Along with orthogonality of the minimizer to the null space, this turns out to be sufficient to show convergence proportional to $\left(\sum_{n=N_{b}+1}^{\infty}\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{2p}d_{n}^{-1}\right)^{\frac{1}{2}}$ of the truncated basis solution to the $\ell^{2}$ problem. The idea here is to show that adding terms to our Fourier basis recovers the solution to the full $\ell^{2}$ problem in the limit $N_{b}\to\infty$.
@@
$\boldsymbol{V}_{N_{b}}\tilde{\boldsymbol{b}}$ $=\boldsymbol{f},$
@@
$\tilde{\boldsymbol{b}}=\boldsymbol{V}_{N_{b}}^{+}\boldsymbol{f},$
@@
This holds for any $\boldsymbol{f}$, so if we define $M_{\tilde{N}}=\left\|\boldsymbol{V}_{N_{b,\text{interp}}}^{+}\right\|_{2}$ , then as long as $N_{b}\geq N_{b,\text{interp}}$,
@@
$\varepsilon_{\text{trunc}}:=\left\|\boldsymbol{f}-\boldsymbol{V}_{N_{b}}\tilde{\boldsymbol{b}}_{\infty}^{\left(N_{b}\right)}\right\|_{2},$
@@
$\boldsymbol{V}_{N_{b}}\left(\tilde{\boldsymbol{b}}_{\infty}^{\left(N_{b}\right)}+\boldsymbol{V}_{N_{b}}^{+}\left(\boldsymbol{f}-\boldsymbol{V}_{N_{b}}\tilde{\boldsymbol{b}}_{\infty}^{\left(N_{b}\right)}\right)\right)=\boldsymbol{f}.$
@@
$\left\|\tilde{\boldsymbol{b}}_{\infty}^{\left(N_{b}\right)}+\boldsymbol{V}_{N_{b}}^{+}\left(\boldsymbol{f}-\boldsymbol{V}_{N_{b}}\tilde{\boldsymbol{b}}_{\infty}^{\left(N_{b}\right)}\right)\right\|_{2}$ $\geq\left\|\tilde{\boldsymbol{b}}\right\|_{2}$
@@
$\varepsilon_{\text{trunc}}^{2}$ $=\sum_{j=1}^{N_{S}}\sum_{k=1}^{\tilde{N}_{j}}\left|f^{\left(j\right)}\left(\boldsymbol{x}^{\left(j\right)}_{k}\right)-\sum_{n=1}^{N_{b}}d_{n}^{-\frac{1}{2}}\tilde{b}_{\infty,n}\mathcal{F}^{\left(j\right)}\phi_{n}\left(\boldsymbol{x}^{\left(j\right)}_{k}\right)\right|^{2}$
$=\sum_{j=1}^{N_{S}}\sum_{k=1}^{\tilde{N}_{j}}\left|\sum_{n=N_{b}+1}^{\infty}d_{n}^{-\frac{1}{2}}\tilde{b}_{\infty,n}\mathcal{F}^{\left(j\right)}\phi_{n}\left(\boldsymbol{x}^{\left(j\right)}_{k}\right)\right|^{2}$
$\leq B_{p}^{2}\tilde{N}\left(\sum_{n=N_{b}+1}^{\infty}\left|\tilde{b}_{\infty,n}\right|^{2}\right)\left(\sum_{n=N_{b}+1}^{\infty}\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{2p}d_{n}^{-1}\right)\text{, by Cauchy-Schwarz},$
where the last line holds for large enough $N_{b}$ by assumption on $\mathcal{F}^{\left(j\right)}$. Finally, from (3.8),
@@
In 1D, an interpolant for a combination of function or derivative interpolation conditions (a Hermite-Birkhoff interpolant) exists under certain conditions when $N_{b}=\tilde{N}$ (Johnson, 1975), but this is not necessarily the case in dimensions greater than one, again due to the Mairhuber-Curtis Theorem (Curtis Jr., 1959; Mairhuber, 1956). An interpolant does, however, always exist with some finite number of basis functions.
In particular, it can be quickly shown that $N_{b,\text{interp}}\leq\left(\left(p+1\right)\tilde{N}_{\text{distinct}}+1\right)^{m}$, where $\tilde{N}_{\text{distinct}}\leq\tilde{N}$ is the number of distinct points in the box $\Omega\subset\mathbb{R}^{m}$. $\tilde{N}_{\text{distinct}}$ can be less than $\tilde{N}$ if one or more points have two or more interpolation conditions imposed. For $m=1$, it has been known for some time (Johnson, 1975, Thm. 1) that we can always construct a trigonometric polynomial with at most $\left(p+1\right)\tilde{N}_{\text{distinct}}+1$ terms that is zero and has all derivatives of up to order $p$ equal to zero on up to $\tilde{N}_{\text{distinct}}$ distinct points of our choosing, except for one chosen point where either the function or one of its derivatives are equal to one. For $m>1$, we can take products of these 1D functions to get a trigonometric polynomial with at most $\left(\left(p+1\right)\tilde{N}_{\text{distinct}}+1\right)^{m}$ terms that is zero and has derivatives equal to zero on a tensor grid of $\tilde{N}_{\text{distinct}}^{m}$ points (that our desired $\tilde{N}_{\text{distinct}}$ points are a subset of), except for one point and derivative of interest. A linear combination of these functions is a feasible interpolant. Experimentally, however, we typically observe $N_{b,\text{interp}}$ to simply be the total number of interpolation conditions ($\tilde{N}$) or not much larger (less than $2\tilde{N}$).
Theorem 2.2 of Chandrasekaran et al. (2018) provides an upper bound on $N_{b,\text{interp}}$ that depends on point separation and will often be less than the simple bound presented above. Note that increasing $N_{b}$ well beyond $N_{b,\text{interp}}$ can improve the error, in accordance with Proposition 9.